Project

General

Profile

Actions

Operations Meeting, in Preparation for Mock Data Run

Pierrick
Steve
Adam Dobbs
Ed Overton
David Adey
Chris Rogers
Yordan Karadzhov

Diffuser
  • IOC is running
  • State machine is near completion, needs testing
  • Run control using the state machine needs to be tested
  • Hardware all okay
TOF
  • HV can't be run on remote
  • HV test on Ckov
  • PMTs all okay
Tracker
  • Cryo - 3+4 are cold (D/S), only these will get wave guide on
  • Cryo - probably cold by Friday
  • Waveguides need to be in; waiting on John Govans to bring them over on forklift
  • Cabling of D/S tracker this afternoon
  • Ground the cryostats
  • Short to be fixed on cryo 3
  • CAEN VME controller needs to be switched to D/S rack, currently they are in the U/S rack
    • D/S ones are in Sheffield
    • Or use SVS; SVS is regarded as a fallback
    • Round of initialisations and simple standalone readout tomorrow
    • Use current Adey C+M to drive DAQ
    • Merge tracker code in with DATE
    • Attempt to readout, using old trigger
    • Attempt to bring in new trigger, need to get to that in next two days
    • Yordan will test Overton FPGA code
    • Use LED system once Alan and Paul arrive on Friday
  • Tracker Controls
    • Merging James's non-expert code; Ed and Pierrick, following TOF HV
    • Hardware things; CAEN controller, moving back to SVS is possible but not tested, CAEN controller is not tested; Pierrick casts doubt on possibility
    • Need to run a new fibre next few days, 3 fibres spare
    • Daresbury needs to finish readout of MKS (vacuum) and temp controllers
    • Need interlock CANBus fixed
  • Unpacking
    • No change
  • OnMon
    • Adey has spoken with Rhys, he will be here later this week to discuss
  • OnRec
    • Dobbs and Adey requested status separately from Melissa, no reply
    • No change
    • Rogers to chase
  • Run plan
    • Need to introduce appropriate delays on LED before it hits the trigger
    • LED needs to pulse first
KL
  • No known problems; it is on SYV527
EMR
  • Two front end boards noisy need replacing
  • Missing power supply for LED
Ckov
  • No known issues
LM
  • Pierrick, when he was installing the IOC trace, he noticed the LM power supply came disconnected; need to check that it has not been damaged
  • Check voltages on PMTs
  • Look for dark currents off the PMT
DAQ
  • No known problems
  • New machines
  • New DATE
  • New linux install
  • General tests, but all looked okay
  • Need to test with real data
Controls
  • EMR control and monitoring of rack is likely to slip
Control room
  • Failed disk on miceserv
  • micethin network broken
  • niggles
  • NAGIOS deployment
  • Network switch okay
  • Reset of a mice thin may break network connection
Datamover
  • Need to update movefiles script; change to Step IV
  • Not expecting automated data mover to run
  • Janusz will be around on 21st January
PPS
  • Nothing new
  • Note that beam stop will be closed for PPS reasons while the target is operating
Beamline
  • Target - S1.9 is now installed, few niggles
    • Will do testing on fibres
    • Swap controller with R78 controller
    • Check quadrature
    • Run up target before Friday lunchtime
    • Need to check with ISIS that there is vacuum
  • Decay solenoid
    • Attempting to fix DS trip on low current this week
    • Currently DS is at 200 A; after today it will be owned by expert
    • Should be in the hall, may be tripping off fridge, need to have easy access to MLCR/Linde fridge PC
  • NC magnets
    • Water from ISIS is marginal, may cause problem
    • Fenced and signs still up in synch, Henry to take signs down later today following tests
  • Beamstop serviced and working
    • Need to be closed for D1 and D2 and target frame lowered
  • Proton absorber
    • No known issues
OnRec
  • Go through with Adam after lunch on Wednesday 14.30 onwards
  • No request for SOC
  • Action on onrec to move onrec02 and 03 into the server

Updated by Rogers, Chris over 8 years ago ยท 1 revisions